# Jacobian Descent For Multi-Objective Optimization

### v1

@@ Abstract
Many optimization problems are inherently multi-objective. To address them, we formalize Jacobian descent (JD), a direct generalization of gradient descent for vector-valued functions. Each step of this algorithm relies on a Jacobian matrix consisting of one gradient per objective. The aggregator, responsible for reducing this matrix into an update vector, characterizes JD. While the multi-task learning literature already contains a variety of aggregators, they often lack some natural properties. In particular, the update should not conflict with any objective and should scale proportionally to the norm of each gradient. We propose a new aggregator specifically designed to satisfy this. Emphasizing conflict between objectives, we then highlight direct applications for our methods. Most notably, we introduce instance-wise risk minimization (IWRM), a learning paradigm in which the loss of each training example is considered a separate objective. On simple image classification tasks, IWRM exhibits promising results compared to the direct minimization of the average loss. The performance of our aggregator in those experiments also corroborates our theoretical findings. Lastly, as speed is the main limitation of JD, we provide a path towards a more efficient implementation.
[…]
@@ 1 Introduction
The field of multi-objective optimization studies minimization of vector-valued objective functions [1, 2, 3, 4]. In deep learning, a widespread approach to train a model with multiple objectives is to combine those into a scalar loss function minimized by stochastic gradient descent. While this method is simple, it comes at the expense of potentially degrading some individual objectives. Without prior knowledge of their relative importance, this is undesirable.
@@
Early works have attempted to extend gradient descent (GD) to consider several objectives simultaneously, and thus several gradients [5, 6]. Essentially, [5, 6] propose some heuristic to prevent the degradation of any individual objective. Several other works have built upon this method, analyzing its convergence properties or extending it to a stochastic setting [7, 8, 9]. Later, this has been applied to multi-task learning to tackle conflict between tasks, illustrated by contradicting gradient directions [10]. Many studies have followed, proposing various other algorithms for the training of multi-task models [11, 12, 13, 14, 15, 16, 17]. They commonly rely on an aggregator that maps a collection of task-specific gradients (a Jacobian matrix) to a shared parameter update.
@@
We propose to unify all such methods under the Jacobian descent (JD) algorithm, specified by an aggregator. [^1] Our library enabling JD with PyTorch is available at https://github.com/TorchJD/torchjd The goal of this algorithm is to minimize a differentiable vector-valued function ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ iteratively without relying on a scalarization of the objective. Besides, we introduce a novel stochastic variant of JD that enables the training of neural networks with a large number of objectives. This unlocks a particularly interesting perspective: considering the minimization of instance-wise loss vectors rather than the usual minimization of the average training loss. As this paradigm is a direct generalization of the well-known empirical risk minimization (ERM) [18], we name it instance-wise risk minimization (IWRM).
With the JD formulation, we can ponder the properties that aggregators should satisfy. The need for a new aggregator emerges from the realization that none of the existing methods is both robust to conflict and proportionally affected by individual gradient magnitudes.
@@
Our contributions are organized as follows: In Section 2, we formalize the JD algorithm and its stochastic variants. We then introduce three critical aggregator properties and define $\mathcal{A}_{\text{UPGrad}}$ to satisfy them. In the smooth convex case, we show convergence of JD with $\mathcal{A}_{\text{UPGrad}}$ to the Pareto front. We present applications for JD and aggregators in Section 3, emphasizing the IWRM paradigm. We then discuss existing aggregators and analyze their properties in Section 4. In Section 5, we report experiments with IWRM optimized with stochastic JD with various aggregators. Lastly, we address computational efficiency in Section 6, giving a path towards an efficient implementation.
@@ 2.1 Jacobian descent
Suppose that ${\boldsymbol{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ is continuously differentiable. Let $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\in\mathbb{R}^{m\times n}$ be the Jacobian matrix of ${\boldsymbol{f}}$ at ${\boldsymbol{x}}$, i.e. $\left[\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\right]_{ij}=\left[\frac{\partial}{\partial x_{j}}{\boldsymbol{f}}({\boldsymbol{x}})\right]_{i}$. If a given ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$ is updated with some ${\boldsymbol{y}}\in{\mathbb{R}}^{n}$, Taylor’s theorem yields
${\boldsymbol{f}}({\boldsymbol{x}}+{\boldsymbol{y}})={\boldsymbol{f}}({\boldsymbol{x}})+\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}+o(\|{\boldsymbol{y}}\|),$ (1)
where $o(\|{\boldsymbol{y}}\|)$ means that $\lim_{\|{\boldsymbol{y}}\|\to 0}\frac{{\boldsymbol{f}}({\boldsymbol{x}}+{\boldsymbol{y}})-{\boldsymbol{f}}({\boldsymbol{x}})-\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}}{\|{\boldsymbol{y}}\|}={\boldsymbol{0}}$. The term ${\boldsymbol{f}}({\boldsymbol{x}})+\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}$ is the first-order Taylor approximation of ${\boldsymbol{f}}({\boldsymbol{x}}+{\boldsymbol{y}})$. Since it depends on ${\boldsymbol{y}}$ only through $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}$, it is sensible to make ${\boldsymbol{y}}$ a function of $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})$. A mapping $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$ reducing such a matrix into a vector is called an aggregator. For any $J\in\mathbb{R}^{m\times n}$, $\mathcal{A}(J)$ is called the aggregation of $J$ by $\mathcal{A}$. To minimize ${\boldsymbol{f}}$, the update is then selected as ${\boldsymbol{y}}=-\eta\mathcal{A}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}))$, where $\eta$ is an appropriate step size.
When $m=1$, the Jacobian consists of a single gradient. In GD, the update is thus simply $-\eta\nabla f({\boldsymbol{x}})$, i.e. the aggregator is the identity. A minimal version of GD is given in Algorithm 1. When $m>1$, the choice of $\mathcal{A}$ is non-trivial and is one of the main subjects of the present work. An elementary version of JD parameterized by some aggregator $\mathcal{A}$ is provided in Algorithm 2.
Input: ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, $0<\eta$, $T\in\mathbb{N}$
@@
       ${\boldsymbol{x}}\leftarrow{\boldsymbol{x}}-\eta\nabla f({\boldsymbol{x}})$
Output: ${\boldsymbol{x}}$
@@
Input: $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$, ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, $0<\eta$, $T\in\mathbb{N}$
for $t\leftarrow 1$ to $T$ do
       ${\boldsymbol{x}}\leftarrow{\boldsymbol{x}}-\eta\mathcal{A}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}))$
Output: ${\boldsymbol{x}}$
Algorithm 2 Jacobian descent with aggregator $\mathcal{A}$
Note that most gradient-based optimization algorithms, e.g. Adam [19], can similarly be extended to the multi-objective setting by substituting the gradient with an aggregation of the Jacobian.
In some settings, the exact computation of the update can be prohibitively slow or even intractable. When dealing with a single objective, the gradient $\nabla f({\boldsymbol{x}})$ can be substituted with an estimation. This is known as stochastic gradient descent (SGD). More generally, stochastic Jacobian descent (SJD) relies on estimates of the aggregation of the Jacobian. We highlight two straightforward methods for this. First, we can compute a stochastic estimation of the Jacobian and aggregate it instead of the true Jacobian. We call this stochastically estimated Jacobian descent (SEJD). Moreover, we can aggregate a matrix whose rows are a random subset of the rows of the true Jacobian. We refer to this as stochastic sub-Jacobian descent (SSJD). This novel approach enables multi-objective optimization with a very large number of objectives.
2.2 Desired properties for aggregators
An inherent challenge of multi-objective optimization is to manage conflicting objectives [10, 11, 12]. Note that substituting the update ${\boldsymbol{y}}=-\eta\mathcal{A}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}))$ into the first-order Taylor approximation ${\boldsymbol{f}}({\boldsymbol{x}})~{}+~{}\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}$ yields ${\boldsymbol{f}}({\boldsymbol{x}})-\eta\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot\mathcal{A}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}))$. In particular, if ${\boldsymbol{0}}\preceq\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot\mathcal{A}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}))$, then no coordinate of the approximation of ${\boldsymbol{f}}$ will increase. A pair of vectors ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathbb{R}}^{n}$ is said to conflict if ${\boldsymbol{x}}^{\top}{\boldsymbol{y}}<0$. Hence, if any row of $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})$ conflicts with $\mathcal{A}(\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}}))$, the corresponding coordinate of ${\boldsymbol{f}}$ will locally increase. When minimizing ${\boldsymbol{f}}$, avoiding conflict between the aggregation and any gradient is thus desirable. In GD, $\nabla{\boldsymbol{f}}({\boldsymbol{x}})$ does indeed not conflict with itself. For aggregators of JD, this motivates the first property.
@@ 2.2 Desired properties for aggregators
Let $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$ be an aggregator. If for all $J\in\mathbb{R}^{m\times n}$, ${\boldsymbol{0}}\preceq J\cdot\mathcal{A}(J)$, then $\mathcal{A}$ is said to be non-conflicting.
For any collection of vectors $C\subseteq{\mathbb{R}}^{n}$, the dual cone of $C$ is $\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}:\forall{\boldsymbol{y}}\in C,0\varleq{\boldsymbol{x}}^{\top}{\boldsymbol{y}}\}$ [20]. Notice that an aggregator $\mathcal{A}$ is non-conflicting if and only if for any $J$, $\mathcal{A}(J)$ is in the dual cone of the rows of $J$.
In a step of GD, the update scales proportionally to the gradient norm. Small gradients will thus lead to small updates, and conversely, large gradients will lead to large updates. Likewise, it would be coherent that the rows of the Jacobian also contribute to the aggregation proportionally to their norm. Scaling each row of a matrix $J\in\mathbb{R}^{m\times n}$ by the corresponding element of a vector ${\boldsymbol{c}}\in{\mathbb{R}}^{m}$ gives $\operatorname*{diag}({\boldsymbol{c}})\cdot J$. This insight can then be formalized as the following property.
@@
Let $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$ be an aggregator. If for all $J\in\mathbb{R}^{m\times n}$, the mapping from any ${\boldsymbol{0}}\prec{\boldsymbol{c}}\in{\mathbb{R}}^{m}$ to $\mathcal{A}(\operatorname*{diag}({\boldsymbol{c}})\cdot J)$ is linear in ${\boldsymbol{c}}$, then $\mathcal{A}$ is said to be linear under scaling.
Finally, (1) highlights that the precision of the first order Taylor approximation ${\boldsymbol{f}}({\boldsymbol{x}})+\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}$ improves as $\|{\boldsymbol{y}}\|$ decreases. For any candidate update ${\boldsymbol{y}}$, its projection ${\boldsymbol{y}}^{\prime}$ onto the row space of $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})$ is such that $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}^{\prime}=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})\cdot{\boldsymbol{y}}$ and $\|{\boldsymbol{y}}^{\prime}\|\varleq\|{\boldsymbol{y}}\|$. This projection thus preserves the value of the approximation while improving its precision. It is therefore sensible to select ${\boldsymbol{y}}$ directly in the span of the rows of $\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})$, i.e. to have a vector of weights ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ such that ${\boldsymbol{y}}=\mathcal{J}\kern-2.0pt{\boldsymbol{f}}({\boldsymbol{x}})^{\top}\cdot{\boldsymbol{w}}$. This holds by design in GD and yields the last desired property for aggregators of JD.
@@
Let $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$ be an aggregator. If for all $J\in\mathbb{R}^{m\times n}$, there exists ${\boldsymbol{w}}\in{\mathbb{R}}^{m}$ such that $\mathcal{A}(J)=J^{\top}\cdot{\boldsymbol{w}}$, then $\mathcal{A}$ is said to be weighted.

### v2

@@ Abstract
Many optimization problems require balancing multiple conflicting objectives. As gradient descent is limited to single-objective optimization, we introduce its direct generalization: Jacobian descent (JD). This algorithm iteratively updates parameters using the Jacobian matrix of a vector-valued objective function, in which each row is the gradient of an individual objective. While several methods to combine gradients already exist in the literature, they are generally hindered when the objectives conflict. In contrast, we propose projecting gradients to fully resolve conflict while ensuring that they preserve an influence proportional to their norm. We prove significantly stronger convergence guarantees with this approach, supported by our empirical results. Our method also enables instance-wise risk minimization (IWRM), a novel learning paradigm in which the loss of each training example is considered a separate objective. Applied to simple image classification tasks, IWRM exhibits promising results compared to the direct minimization of the average loss. Additionally, we outline an efficient implementation of JD using the Gramian of the Jacobian matrix to reduce time and memory requirements.
@@ 1 Introduction
The field of multi-objective optimization studies minimization of vector-valued objective functions (Sawaragi et al., 1985; Ehrgott, 2005; Branke, 2008; Deb et al., 2016). In deep learning, a widespread approach to train a model with multiple objectives is to combine those into a scalar loss function minimized by stochastic gradient descent. While this method is simple, it comes at the expense of potentially degrading some individual objectives. Without prior knowledge of their relative importance, this is undesirable.
@@
Early works have attempted to extend gradient descent (GD) to consider several objectives simultaneously, and thus several gradients (Fliege & Svaiter, 2000; Désidéri, 2012). Essentially, they propose a heuristic to prevent the degradation of any individual objective. Several other works have built upon this method, analyzing its convergence properties or extending it to a stochastic setting (Fliege et al., 2019; Poirion et al., 2017; Mercier et al., 2018). Later, this has been applied to multi-task learning to tackle conflict between tasks, illustrated by contradicting gradient directions (Sener & Koltun, 2018). Many studies have followed, proposing various other algorithms for the training of multi-task models (Yu et al., 2020; Liu et al., 2021a, b; Lin et al., 2021; Navon et al., 2022; Senushkin et al., 2023; Chen et al., 2020). They commonly rely on an aggregator that maps a collection of task-specific gradients (a Jacobian matrix) to a shared parameter update.
We propose to unify all such methods under the Jacobian descent (JD) algorithm, specified by an aggregator. [^1] Our library enabling JD with PyTorch is available at https://github.com/TorchJD/torchjd This algorithm aims to minimize a differentiable vector-valued function ${\bm{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ iteratively without relying on a scalarization of the objective. Under this formulation, the existing methods are simply distinguished by their aggregator. Consequently, studying its properties is essential for understanding the behavior and convergence of JD. Under significant conflict, existing aggregators often fail to provide strong convergence guarantees. To address this, we propose $\mathcal{A}_{\text{UPGrad}}$, specifically designed to resolve conflicts while naturally preserving the relative influence of individual gradients.
@@
Furthermore, we introduce a novel stochastic variant of JD that enables the training of neural networks with a large number of objectives. This unlocks a particularly interesting perspective: considering the minimization of instance-wise loss vectors rather than the usual minimization of the average training loss. As this paradigm is a direct generalization of the well-known empirical risk minimization (ERM) (Vapnik, 1995), we name it instance-wise risk minimization (IWRM).
@@
Our contributions are organized as follows: In Section 2, we formalize the JD algorithm and its stochastic variants. We then introduce three important aggregator properties and define $\mathcal{A}_{\text{UPGrad}}$ to satisfy them. In the smooth convex case, we show convergence of JD with $\mathcal{A}_{\text{UPGrad}}$ to the Pareto front. We present applications for JD and aggregators in Section 3, emphasizing the IWRM paradigm. We then discuss existing aggregators and analyze their properties in Section 4. In Section 5, we report experiments with IWRM optimized with stochastic JD with various aggregators. Lastly, we address computational efficiency in Section 6, giving a path towards an efficient implementation.
@@ 2.1 Jacobian descent
Suppose that ${\bm{f}}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{m}$ is continuously differentiable. Let $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\in\mathbb{R}^{m\times n}$ be the Jacobian matrix of ${\bm{f}}$ at ${\bm{x}}$, i.e.
$\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})=\begin{bmatrix}\nabla f_{1}({\bm{x}})^{\top}\\ \nabla f_{2}({\bm{x}})^{\top}\\ \vdots\\ \nabla f_{m}({\bm{x}})^{\top}\end{bmatrix}=\begin{bmatrix}\frac{\partial}{\partial x_{1}}f_{1}({\bm{x}})&\frac{\partial}{\partial x_{2}}f_{1}({\bm{x}})&\cdots&\frac{\partial}{\partial x_{n}}f_{1}({\bm{x}})\\ \frac{\partial}{\partial x_{1}}f_{2}({\bm{x}})&\frac{\partial}{\partial x_{2}}f_{2}({\bm{x}})&\cdots&\frac{\partial}{\partial x_{n}}f_{2}({\bm{x}})\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial}{\partial x_{1}}f_{m}({\bm{x}})&\frac{\partial}{\partial x_{2}}f_{m}({\bm{x}})&\cdots&\frac{\partial}{\partial x_{n}}f_{m}({\bm{x}})\\ \end{bmatrix}$ (1)
Given ${\bm{x}},{\bm{y}}\in{\mathbb{R}}^{n}$, Taylor’s theorem yields
${\bm{f}}({\bm{x}}+{\bm{y}})={\bm{f}}({\bm{x}})+\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot{\bm{y}}+o(\|{\bm{y}}\|),$ (2)
where $o(\|{\bm{y}}\|)$ indicates that $\lim_{\|{\bm{y}}\|\to 0}\frac{{\bm{f}}({\bm{x}}+{\bm{y}})-{\bm{f}}({\bm{x}})-\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot{\bm{y}}}{\|{\bm{y}}\|}={\bm{0}}$. The term ${\bm{f}}({\bm{x}})+\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot{\bm{y}}$ is the first-order Taylor approximation of ${\bm{f}}({\bm{x}}+{\bm{y}})$. Via this approximation, we aim to select a small update ${\bm{y}}$ that reduces ${\bm{f}}({\bm{x}}+{\bm{y}})$, ideally achieving ${\bm{f}}({\bm{x}}+{\bm{y}})\preceq{\bm{f}}({\bm{x}})$. As the approximation depends on ${\bm{y}}$ only through $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot{\bm{y}}$, selecting the update based on the Jacobian is natural. A mapping $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$ reducing such a matrix into a vector is called an aggregator. For any $J\in\mathbb{R}^{m\times n}$, $\mathcal{A}(J)$ is called the aggregation of $J$ by $\mathcal{A}$.
To minimize ${\bm{f}}$, consider the update ${\bm{y}}=-\eta\mathcal{A}\big{(}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\big{)}$, where $\eta$ is an appropriate step size, and $\mathcal{A}$ is an appropriate aggregator. Jacobian descent simply consists in applying this update iteratively, as shown in Algorithm 1. To put it into perspective, we also provide a minimal version of GD in Algorithm 2. Remarkably, when $m=1$, the Jacobian has a single row, so GD is a special case of JD where the aggregator is the identity.
Input: ${\bm{x}}\in{\mathbb{R}}^{n}$, $0<\eta$, $T\in\mathbb{N}$, $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$
@@
       ${\bm{x}}\leftarrow{\bm{x}}-\eta\mathcal{A}\big{(}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\big{)}$
Output: ${\bm{x}}$
Algorithm 1 Jacobian descent with aggregator $\mathcal{A}$
Input: ${\bm{x}}\in{\mathbb{R}}^{n}$, $0<\eta$, $T\in\mathbb{N}$
for $t\leftarrow 1$ to $T$ do
       ${\bm{x}}\leftarrow{\bm{x}}-\eta\nabla f({\bm{x}})$
Output: ${\bm{x}}$
@@
Note that other gradient-based optimization algorithms, e.g. Adam (Kingma & Ba, 2014), can similarly be extended to the multi-objective case.
In some settings, the exact computation of the update can be prohibitively slow or even intractable. When dealing with a single objective, stochastic gradient descent (SGD) replaces the gradient $\nabla f({\bm{x}})$ with some estimation. More generally, stochastic Jacobian descent (SJD) relies on estimates of the aggregation of the Jacobian. One approach, that we call stochastically estimated Jacobian descent (SEJD), is to compute and aggregate an estimation of the Jacobian. Alternatively, when the number of objectives is very large, we propose to aggregate a matrix whose rows are a random subset of the rows of the true Jacobian. We call this approach stochastic sub-Jacobian descent (SSJD).
2.2 Desirable properties for aggregators
An inherent challenge of multi-objective optimization is to manage conflicting objectives (Sener & Koltun, 2018; Yu et al., 2020; Liu et al., 2021a). Substituting the update ${\bm{y}}=-\eta\mathcal{A}\big{(}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\big{)}$ into the first-order Taylor approximation ${\bm{f}}({\bm{x}})~{}+~{}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot{\bm{y}}$ yields ${\bm{f}}({\bm{x}})-\eta\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot\mathcal{A}\big{(}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\big{)}$. In particular, if ${\bm{0}}\preceq\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot\mathcal{A}\big{(}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\big{)}$, then no coordinate of the approximation of ${\bm{f}}$ will increase. A pair of vectors ${\bm{x}},{\bm{y}}\in{\mathbb{R}}^{n}$ is said to conflict if ${\bm{x}}^{\top}{\bm{y}}<0$. Hence, for a sufficiently small $\eta$, if any row of $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})$ conflicts with $\mathcal{A}\big{(}\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\big{)}$, the corresponding coordinate of ${\bm{f}}$ will increase. When minimizing ${\bm{f}}$, avoiding conflict between the aggregation and any gradient is thus desirable, motivating the first property.
@@ 2.2 Desirable properties for aggregators
Let $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$ be an aggregator. If for all $J\in\mathbb{R}^{m\times n}$, ${\bm{0}}\preceq J\cdot\mathcal{A}(J)$, then $\mathcal{A}$ is said to be non-conflicting.
For any collection of vectors $C\subseteq{\mathbb{R}}^{n}$, the dual cone of $C$ is $\{{\bm{x}}\in{\mathbb{R}}^{n}:\forall{\bm{y}}\in C,0\varleq{\bm{x}}^{\top}{\bm{y}}\}$ (Boyd & Vandenberghe, 2004). Notice that an aggregator $\mathcal{A}$ is non-conflicting if and only if for any $J$, $\mathcal{A}(J)$ is in the dual cone of the rows of $J$.
In a step of GD, the update scales proportionally to the gradient norm. Small gradients thus lead to small updates, and conversely, large gradients lead to large updates. To maintain coherence with GD, it would be natural that the rows of the Jacobian also contribute to the aggregation proportionally to their norm. Scaling each row of $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})$ by the corresponding element of some vector ${\bm{c}}\in{\mathbb{R}}^{m}$ yields $\operatorname*{diag}({\bm{c}})\cdot\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})$. This insight can then be formalized as the following property.
@@
Let $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$ be an aggregator. If for all $J\in\mathbb{R}^{m\times n}$, the mapping from any ${\bm{0}}\prec{\bm{c}}\in{\mathbb{R}}^{m}$ to $\mathcal{A}\big{(}\operatorname*{diag}({\bm{c}})\cdot J\big{)}$ is linear in ${\bm{c}}$, then $\mathcal{A}$ is said to be linear under scaling.
Finally, as $\|{\bm{y}}\|$ decreases asymptotically to 00, the precision of the first-order Taylor approximation ${\bm{f}}({\bm{x}})+\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot{\bm{y}}$ improves, as highlighted in (2). The projection ${\bm{y}}^{\prime}$ of any candidate update ${\bm{y}}$ onto the span of the rows of $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})$ satisfies $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot{\bm{y}}^{\prime}=\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})\cdot{\bm{y}}$ and $\|{\bm{y}}^{\prime}\|\varleq\|{\bm{y}}\|$, so this projection decreases the norm of the update while preserving the value of the approximation. Without additional information about ${\bm{f}}$, it is thus reasonable to select ${\bm{y}}$ directly in the row span of $\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})$, i.e. to have a vector of weights ${\bm{w}}\in{\mathbb{R}}^{m}$ satisfying ${\bm{y}}=\mathcal{J}\kern-2.0pt{\bm{f}}({\bm{x}})^{\top}\cdot{\bm{w}}$. This yields the last desirable property.
@@
Let $\mathcal{A}:\mathbb{R}^{m\times n}\to{\mathbb{R}}^{n}$ be an aggregator. If for all $J\in\mathbb{R}^{m\times n}$, there exists ${\bm{w}}\in{\mathbb{R}}^{m}$ satisfying $\mathcal{A}(J)=J^{\top}\cdot{\bm{w}}$, then $\mathcal{A}$ is said to be weighted.